OpenStudy (anonymous):

evaluate (11^1/9)-^12

OpenStudy (amistre64):

flip it; multiply exps; simplify to 4/3 and cbrt the denominator of 11^4

OpenStudy (amistre64):

\[\frac{1}{\sqrt[3]{11^4}}\]

OpenStudy (anonymous):

can 11 be cubed?

OpenStudy (amistre64):

not that I can tell; nothing perfect cubes to 11

OpenStudy (amistre64):

at least no integers do

OpenStudy (anonymous):

so if it cant be cubed I just work out 11^4?

OpenStudy (amistre64):

well, notice that 11^4 = 11^3 * 11 \[\frac1{11\sqrt[3]{11}}\] \[\frac1{11\sqrt[3]{11}}*\frac{\sqrt[3]{11^2}}{\sqrt[3]{11^2}}\] \[\frac{\sqrt[3]{11^2}}{121}\] maybe

OpenStudy (anonymous):

my book gives the answer 3/10

OpenStudy (amistre64):

that seems a bit contrived; you sure your matching up the right problem/answer ?

OpenStudy (anonymous):

yes doubled checked before posting

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=11^%28-4%2F3%29

OpenStudy (anonymous):

sorry, i question is incorrect

OpenStudy (anonymous):

evaluate (11^1/9)-^1/2

OpenStudy (amistre64):

still not 3/10 http://www.wolframalpha.com/input/?i=%2811^%281%2F9%29%29^%28-1%2F2%29

OpenStudy (anonymous):

could my textbook be wrong?

OpenStudy (amistre64):

its possible, but there tends to be a myriad of reasons as to what might be going awry

OpenStudy (amistre64):

B^a^b = B^ab is the basic rule here; then simplify the results

OpenStudy (radar):

\[(11^{1/9})^{-1/2}=11^{-1/18}=1/\sqrt[18]{11}\]

OpenStudy (anonymous):

@ radar my textbook gives the answer 3/10

OpenStudy (anonymous):

@ Hollywood_chrissy Would it be possible to post a digital picture of the expression shown in your textbook?

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

Question 41.

OpenStudy (anonymous):

oooooooooh is this \[(11\tfrac{1}{9})^{-\frac{1}{2}}\]

OpenStudy (anonymous):

then get rid of that stupid improper fraction and write \[(\frac{100}{9})^{-\frac{1}{2}}=(\frac{9}{100})^{\frac{1}{2}}=\sqrt{\frac{100}{9}}=\frac{10}{3}\]

OpenStudy (anonymous):

lol, brilliant, thank you

OpenStudy (anonymous):

yw

OpenStudy (amistre64):

see, that was one myraidecimal circumstance lol

OpenStudy (anonymous):

sorry??

OpenStudy (radar):

Thanks for clearing that up!